Please help me solve this. Let f(x)=3x^2-4 and g(x)=x^2-3x-4. Find each of the following. A) (f+g)(2k) B)(f/g)(-1)
Answer by stanbon(75887)

Please help me solve this. Let f(x)=3x^2-4 and g(x)=x^2-3x-4. Find each of the following.
-------------------------------
A) (f+g)(2k) = f(2k) + g(2k) = [3(2k)^2-4] + [(2k)^2-3(2k)-4]
---
= 12k^2 - 4 + 4k^2 - 6k -4
----
= 16k^2 -6k - 8
========================

B)(f/g)(-1) = f(-1)/g(-1) = [3(-1)^2-4]/[(-1)^2 -3(-1)-4]
-------
= [3-4]/[1+4-4
-------
= -1/1
----
= -1
---------------------
Cheers,
Stan H.
---------